OpenStudy (anonymous):

solve the p using the addition principle -(3)/9 (17) + p= (7)/(12) p=

OpenStudy (anonymous):

-1/51 +p =7/12 51p-1=7/12 * 51 51p-1=7/4 * 17 51p-1=119/4 51p=119/4 +1 51p=123/4 p=123/204